OpenStudy (anonymous):

24x^2-46x=18 solve the equation check your solution show work

OpenStudy (anonymous):

Okay same thing: 24x^2-46x-18=0 factor 2 (3 x+1) (4 x-9) x = -1/3 x = 9/4

OpenStudy (anonymous):

As before, you can download our software for use during your math class, to check homework, and during the PSAT test (since you are a freshman).

OpenStudy (anonymous):

ill think about it